OpenStudy (anonymous):

Solve equation: n^3-121n=0

OpenStudy (callisto):

n^3-121n=0 n(n^2 -121)=0 n ( n+11)(n-11) =0 n=0 or n+11 =0 or n-11=0 n=-11 or n =11